Exercise 9-20 (Part Level Submission) Crane Company purchased a new machine on October 1, 2017, at a cost of $88,700. The company estimated that the machine has a salvage value of $7,400. The machine is expected to be used for 63,100 working hours during its 8-year life. Compute depreciation using the following methods in the year indicated. Collapse question part (a) Incorrect answer. Your answer is incorrect. Try again. Declining-balance using double the straight-line rate for 2017 and 2018. (Round answers to 0 decimal places, e.g. 125) 2017 2018 Depreciation using the Declining-balance method $Entry field with incorrect answer $Entry field with incorrect answer
Answer #1
Double Declining-balance rate 25% =1/8*2
Depreciation using the Declining-balance method
2017 5544 =88700*25%*3/12
2018 20789 =(88700-5544)*25%